OpenStudy (anonymous):

If the sum of p, q, r, and s is 16, what is the avergae of p+2s, r-q, 2q+5, and 7-s?

OpenStudy (anonymous):

average*

OpenStudy (klimenkov):

Average of 4 values \(a,b,c,d\) is \(\frac{a+b+c+d}4\). Try to use this.

OpenStudy (anonymous):

but how do i relate that to the average of the four terms in the last part

OpenStudy (klimenkov):

Write the average of your values right here.

OpenStudy (anonymous):

4

OpenStudy (klimenkov):

Average of what values did you write?

OpenStudy (anonymous):

(p+2s)+(r-q)+(2q+5)+(7-s)/4

OpenStudy (klimenkov):

So why did you get 4?

OpenStudy (anonymous):

i thought u asked for the average of p q r and s

OpenStudy (klimenkov):

What is the problem for you now?

OpenStudy (anonymous):

how to find the average of that using 16

OpenStudy (anonymous):

because q+ r+ p+ and s = 16

OpenStudy (klimenkov):

Try to simplify this ((p+2s)+(r-q)+(2q+5)+(7-s))/4.

OpenStudy (anonymous):

\[p+2s+r-q+2q+5+ 7-s=12+p+q+r+s=12+16=28 \]\[\frac{28}{4}=7 \]

OpenStudy (anonymous):

p+s+r+q+7/4

OpenStudy (klimenkov):

Do you see the last step to get the answer?

OpenStudy (anonymous):

add 16 and 7 and divide by 4?

OpenStudy (anonymous):

i mean 16 and 12

OpenStudy (anonymous):

so 28/4=7

OpenStudy (klimenkov):

I think you have made a mistake.

OpenStudy (anonymous):

thx i got it!

OpenStudy (klimenkov):

Yes. You are right!

OpenStudy (anonymous):

:D

OpenStudy (klimenkov):

Ask me anything you like. You are welcome.